### v2

# Erasing Conceptual Knowledge from Language Models

Abstract
In this work, we propose Erasure of Language Memory (ELM), an approach to concept-level unlearning built on the principle of matching the distribution defined by an introspective classifier. Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts. ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative analysis shows that ELM achieves superior performance across key metrics, including near-random scores on erased topic assessments, maintained coherence in text generation, preserved accuracy on unrelated benchmarks, and robustness under adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
1 Introduction
What does it mean for a language model to ”unlearn” a concept? While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about biological weapons rather than just a few training examples containing that information. In this paper we examine how concept-level unlearning leads to a new approach to knowledge removal in language models.
Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to ”un-train” specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content. Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts. These approaches lack a principled objective defining successful concept erasure. They focus on technical mechanisms like reversing gradients, altering training data, or randomizing activations without a clear target for the model’s modified behavior.
We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge. Our key insight is that language models can act as their own critics: they can evaluate whether a piece of text demonstrates knowledge of a particular concept. This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification. Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan & Russinovich, 2023) which modifies training data but fails to fully eliminate concept knowledge, ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
Our method achieves this through targeted fine-tuning that reduces the likelihood of generating content the model identifies as related to the concept being erased. When combined with low-rank adaptation of specific layers, this approach effectively eliminates concept knowledge while preserving general capabilities. The self-classification framework also provides a natural way to ensure the model maintains coherent text generation even when prompted about erased concepts—it learns to generate alternative content that it would not classify as demonstrating the target knowledge.
Prior unlearning methods face fundamental limitations when evaluated holistically. Our proposed evaluation desiderata—innocence (lack of target knowledge), specificity (preserved capabilities), and seamlessness (coherent generation)—reveal these tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks, as well as literary domain erasure, we measure the effectiveness of ELM’s principled approach to concept removal compared to previous methods. Our evaluation using both standard metrics and adversarial probing shows that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
2 Related work
Machine Unlearning
The idea of removing specific data from machine learning models, known as machine unlearning, has gained attention in recent years, initially motivated by privacy concerns (Cao & Yang, 2015; Harding et al., 2019). Early methods focused on efficiently removing individual training examples or facts from models (Golatkar et al., 2020; Ma et al., 2022; Jang et al., 2022a). However, most existing benchmarks evaluate unlearning on artificially created deletion sets (Choi & Na, 2023; Goel et al., 2022; Maini et al., 2024), in contrast to our focus on real-world distributions of broad conceptual knowledge.
Erasing broad conceptual knowledge from LLMs
New approaches to machine unlearning have recently gained traction on the problem of removing dangerous capabilities from LLMs (Lynch et al., 2024; Ilharco et al., 2023; Jang et al., 2022b; Lu et al., 2022; Yu et al., 2023; Casper et al., 2024; Eldan & Russinovich, 2023). Our work is most directly comparable to three state-of-the-art techniques: Representation Misdirection for Unlearning (RMU) (Li et al., 2024), WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023), and Representation Noising (RepNoise) (Rosati et al., 2024). RMU fine-tunes models to align internal activations with random scaled vectors when processing targeted concepts. WHP (Eldan & Russinovich, 2023) employs a two-stage approach, first training a reinforced model for the concept being erased and then training an unlearned model to behave differently on the reinforced logits. RepNoise (Rosati et al., 2024) focuses on removing information about harmful representations across all layers of the LLM, by doing gradient ascent along with some representation noising (training internal representations to match Gaussian noise). While all these methods successfully reduce model performance on erased knowledge, our measurements reveal that these previous approaches fall short in one or more of the three erasing goals. In contrast, our current work erases concepts by fine-tuning towards a principled target distribution that is designed to balance all three desiderata of innocence, specificity, and seamlessness.
Other methods such as LLMU (Yao et al., 2023), SSD (Foster et al., 2024), and SCRUB (Kurmanji et al., 2024) have also been proposed but face significant limitations. LLMU uses a differentiable membership classifier to identify unlearning set examples but struggles with imprecisely defined target distributions (see Li et al., 2024). SSD offers a retraining-free method for removing specific samples but is limited to individual examples rather than broader knowledge domains. SCRUB aims to remove sensitive data while preserving model performance on other tasks but requires access to the full training dataset, which is often impractical. Importantly, these methods were compared by RMU (Li et al., 2024) and found to be less effective for erasing broad conceptual knowledge.
Distilling generative model outputs.
Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior. For example Askell et al. (2021) and Bai et al. (2022) prompt an unsafe model into behaving more safely, and then distil the prompted model. Similarly, Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. (Rosati et al., 2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens. Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods. Prior works like Emulator and DeRa leverage probability ratios for behavioral modification - Emulator (Mitchell et al., 2023) scales fine-tuning changes between model sizes, while DeRa (Liu et al., 2024) controls alignment strength during inference. In contrast, ELM introduces a simpler, principled approach focused specifically on reducing the likelihood of knowledge concept generation.
Erasing in generative image models
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023). We take inspiration from (Gandikota et al., 2023) to reduce the likelihood of a concept being generated.
3 Background
3.1 Transformer
We focus on autoregressive transformer-based models (Tunstall et al., 2023a; Dubey et al., 2024). Given an input token sequence , these models output a probability distribution over a vocabulary to predict the next token:
The model architecture comprises stacked blocks, each usually containing a multi-head attention (MHA) layer followed by a multi-layer perceptron (MLP). These layers are stacked on top of each other, and their outputs are added to the residual stream of the model. MHA layers play an important role in bringing information from previous input tokens, while MLP layers are found to play the role of knowledge banks (Meng et al., 2022; Geva et al., 2023). These calculations can be summarized as follows,
where denotes the hidden state for token at layer .
Previous research (Meng et al., 2022) has localized much model knowledge within early to mid-layer blocks. As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
3.2 Transformer Next Token Prediction: A Classification Perspective
Language models are typically viewed through the lens of autoregressive sequence modeling, where the model predicts the next token given a sequence of previous tokens. However, these models can also be understood as powerful classifiers. To illustrate this connection, let’s examine both perspectives:
Traditional Perspective: The standard way to describe an autoregressive language model is through the factorization:
where the model predicts future tokens conditioned on previous tokens .
Classification Perspective: We can reformulate this using Bayes’ rule to obtain:
In this view, the model acts as a classifier where previous tokens serve as class labels for the text that follows. For instance, if we provide the context as ”The following text is about general biology”, the term represents the probability that the generated text belongs to the ”general biology” class.
This classification perspective enables us to manipulate the model’s knowledge by adjusting these classification probabilities. We can define a target distribution that modifies how strongly the model associates generated text with specific concepts:
where is a scaling factor that controls the strength of the modification. When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
To implement this in practice with an autoregressive language model, we apply Bayes’ rule again:
In this form, both the numerator and denominator can be read out of the pretrained language model. Eq 6 is particularly powerful for concept unlearning because it allows us to use the language model itself as both the generator and classifier of content. For example, to unlearn bioweapons expertise, we can reduce the model’s probability of generating text that would naturally follow the context ”You are an expert in bioweapons.” This approach requires no external classifier - we leverage the model’s own understanding of concept relationships as encoded in its next-token predictions. Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
3.3 Parameter-Efficient Fine-tuning via Low-Rank Adapters
Fine-tuning large pre-trained language models involves large sets of parameters, which can result in overfitting. To address this issue, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021), a method that introduces a factorization of weights to significantly reduce the number of parameters. Consider a pre-trained model layer with weights , where and denote input and output dimensions, respectively. LoRA decomposes the weight update as follows:
Here, represents a small rank (), constraining the update to a low-dimensional subspace. By optimizing only and while keeping fixed, LoRA significantly reduces the trainable parameters. During inference, the update can be efficiently incorporated into the weights:
The LoRA approach offers a balance between adaptation capability, reduced overfitting, and parameter efficiency, making it particularly suitable for our setting of fine-tuning large-scale models.
4 Method
We introduce Erasure of Language Memory (ELM), an approach that reformulates concept unlearning through introspective classification. While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge.
Specifically, ELM uses the language model itself as its own introspective classifier, as established in Section 3.2. Rather than training external classifiers or explicitly manipulating representations, we leverage the model’s next-token predictions to identify and reduce the likelihood of generating concept-specific content. This allows us to operate directly on the model’s probability distributions, providing a principled approach to unlearning.
The core of our method is a self-classification objective that reduces the likelihood of generating text that the model would classify as containing the target concept. Given a dataset containing text related to the target concept, we modify the model’s predicted probabilities to diverge from generating content related to the concept being erased and instead favor generating content related to a safer alternative :
where controls the strength of knowledge modification. is the concept to be erased, and is an alternative concept. The corresponding loss is:
In practice, we encounter two main challenges when implementing this objective. First, knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts. To address this, we preserve the model’s behavior on a set of related but safe concepts by matching its original distribution on retain data, , containing unrelated text:
Second, the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts. We address this by applying our core objective (Equation 4) during inference to generate synthetic training examples, then using these to train the model in an autoregressive setting to maintain coherent generation:
The final training objective combines these terms with appropriate weights:
We implement this through low-rank adapters attached to early model layers, allowing precise modification of the model’s knowledge while maintaining its broader capabilities. This approach provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
5 Experiments
5.1 Experimental Setup
Benchmarks.
Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset (Li et al., 2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs). To demonstrate ELM’s versatility, we also employ a modified version of the Harry Potter MCQ dataset (Lynch et al., 2024), expanded from binary to quaternary choices for consistency with other benchmarks. This diverse set of tasks allows us to assess ELM’s erasure effectiveness across different domains and knowledge types.
Models.
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-8B, and Llama3-8B-instruct (Dubey et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from (Eldan & Russinovich, 2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
Baselines.
Data.
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 4), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by (Qi et al., 2024).
Evaluation Metrics.
We assess our method, Erasure of Language Memory (ELM), along four key dimensions:
1.
Innocence: We employ multiple-choice questions (MCQs) related to the target erased class to evaluate contextual knowledge extraction. Additionally, we analyze probing accuracies across internal model layers to detect any traces of latent knowledge.
2.
Seamlessness: To measure the model’s ability to generate fluent text when prompted with erased concepts, we assess the reverse perplexity of generated samples on forget set prompts using an independent language model. We generate text from edited models and run it through a different base model, measuring the perplexity of the text as per the second model (R-PPL). This approach quantifies fluency without relying on potentially biased self-perplexity scores.
3.
Specificity: We evaluate the modified model on standard benchmarks unrelated to the erased content to ensure that the erasure process does not degrade overall model performance.
4.
Robustness: We test against adversarial attacks like GCG (Zou et al., 2023) to understand the model’s tendency to display concept knowledge post-erasure.
5.2 Erasing WMDP Concepts
We evaluate ELM’s performance on erasing biosecurity and cybersecurity concepts from the Weapons of Mass Destruction Proxy (WMDP) dataset (Li et al., 2024). Table 1 presents a comprehensive comparison of ELM against baseline methods RMU (Li et al., 2024) and RepNoise (Rosati et al., 2024) across multiple models and benchmarks.
| Model | Method | Innocence () | Specificity () | Seamlessness | ||
| Bio | Cyber | MMLU | MT-Bench | R-PPL () | ||
| Zephyr-7B | Original | 64.4 | 44.3 | 58.5 | 7.3 | 6.0 |
| RMU | 30.5 | 27.3 | 57.5 | 7.2 | 24.8 | |
| RepNoise | 29.7 | 37.7 | 53.3 | 6.6 | 25.0 | |
| Ours | 29.7 | 27.2 | 56.6 | 7.1 | 10.9 | |
[…]
| RMU | 33.5 | 28.7 | 27.1 | 1.0 | 29.9 | |
| RepNoise | 35.3 | 39.6 | 55.0 | 2.1 | 26.7 | |
| Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
[…]
| RMU | 66.8 | 45.8 | 56.5 | 7.4 | 3.0 | |
| RepNoise | 59.9 | 44.1 | 60.1 | 6.7 | 3.5 | |
| Ours | 32.2 | 27.2 | 61.6 | 7.7 | 7.4 | |
[…]
| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
As shown in Table 1, ELM consistently achieves near-random performance (random guess is 25%) on erased WMDP concepts (Bio and Cyber) while maintaining high scores on general knowledge (MMLU) and language understanding (MT-Bench) tasks. Notably, ELM demonstrates superior fluency when generating text related to erased concepts, as evidenced by lower reverse perplexity scores compared to RMU and RepNoise.
5.3 Ablation Study
We conduct ablation experiments to analyze the contribution of each loss component in ELM. Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms.
First, proves crucial for innocence. Removing significantly degrades erasure performance, with WMDP scores remaining close to the original model. While randomizing logits as a proxy for erasure (replacing in Eq. 10 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs.
The retain term (Eq. 11) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept. Replacing it with random text from WikiText (Merity et al., 2016) slightly reduces fluency while maintaining erasure effectiveness. Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance.
| Setup | Innocence () | Specificity () | Seamlessness () | ||
| Bio | Cyber | MMLU | MT-Bench | RPPL | |
| w/o | 64.8 | 42.7 | 58.0 | 6.9 | 2.7 |
| w/o | 24.3 | 25.8 | 23.6 | 1.2 | 22.0 |
| w/o | 27.6 | 26.4 | 55.7 | 6.6 | 29.8 |
[…]
| Random Fluency | 29.8 | 30.0 | 56.6 | 6.6 | 13.1 |
| Random Both | 51.3 | 30.6 | 58.4 | 6.7 | 9.54 |
[…]
The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings below:
5.4 Specificity Analysis
To assess the specificity of our erasure method, we examine its impact on related MMLU classes. Figure 2 shows the performance of ELM and RMU on related safe concepts whose accuracies has to remain high (higher is better) when WMDP-bio and WMDP-cyber knowledge is erased. We find that both the methods reduces the accuracies slightly on closely-related safer concepts.
5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) (Zou et al., 2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below. We provide GCG attack details in Appendix C. We also show ELM’s robustness to more attacks in Appendix H.
5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by (Li et al., 2024).
The results in Figure 3(a) reveal distinct knowledge retention patterns across methods. ELM and RMU achieve effective erasure, maintaining low probe accuracies across all layers for both biosecurity and cybersecurity MCQs. In contrast, RepNoise shows partial retention, particularly for cybersecurity concepts.
Analysis of activation norms, in Figure 3(b), further highlights the differences. Both ELM and RMU induce out-of-distribution activations in early layers for the forget set, but while RMU continues to exhibits persistent activation norm disruption across all layers, ELM activation norms return to baseline behavior in middle layers. This suggests altered initial processing of erased concepts during knowlege retrieval while preserving text-prediction behavior in later stages. We hypothesize that the late-layer activation norm disruption in RMU impacting overall model fluency. RepNoise shows minimal changes in activation norms, consistent with its less aggressive erasure approach.
5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to (Lynch et al., 2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge. RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
| Method | Innocence () | Specificity () | Seamless () |
| HP-MCQ | MMLU | R-PPL | |
| Original | 66.4 | 47.0 | 3.6 |
| RMU | 51.0 | 44.2 | 3.7 |
| WHP | 58.6 | 43.1 | 3.4 |
| ELM | 38.3 | 45.3 | 3.4 |
6 limitations
While ELM effectively removes targeted concepts through introspective classification, several limitations merit investigation. The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge. Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence - suggesting that our probability modification may be overly aggressive in some cases. The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base. Further work is needed to develop more granular techniques for selective knowledge modification while preserving complex conceptual inter-dependencies.
7 Conclusion
This work reframes the challenge of machine unlearning for large language models, shifting from traditional sample-based approaches to concept-oriented unlearning through introspective classification. Our proposed Erasure of Language Memory (ELM) method demonstrates that effective concept unlearning requires modifying the model’s output distribution based on its own ability to recognize and evaluate knowledge. By using low-rank model updates guided by the model’s introspective classification, ELM achieves targeted concept removal while preserving the model’s broader capabilities. Our experiments show that this approach overcomes limitations of previous methods like gradient ascent or representation disruption, as evidenced by near-random performance on multiple-choice questions related to erased concepts while maintaining accuracy on other tasks. Furthermore, ELM’s resistance to adversarial attacks validates our hypothesis that concept unlearning should leverage the model’s own understanding of its knowledge. In addition to providing a practical solution for concept erasure, we have established a foundation for more comprehensive evaluation of knowledge erasure in language models.
Acknowledgements
RG, and DB are supported by Open Philanthropy and National Science Foundation (Grant Number: NSF-2403303). SF is supported by Open Phil and Khoury Distinguished Fellowship. SM participated in this work while a postdoctoral researcher at Northeastern University supported by an Open Philanthropy grant.
Code
ELM is available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
Appendix A Details on metrics
Multiple Choice Questions.
To measure the multiple choice question accuracy across the different models and erasure methods, we use the lm-evaluation-harness library by EleutherAI (Gao et al., 2024).
MT-Bench.
We employ the single evaluation mode on MT-Bench, using gpt-4o-2024-05-13 as the judge.
Reverse Perplexity (R-PPL).
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP (Li et al., 2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
Appendix B Baseline Methods
B.1 WMDP Results
RMU (Li et al., 2024).
We directly download the best Zephyr-7B RMU model from the WMDP authors (https://huggingface.co/cais/Zephyr_RMU) for testing. For Mistral, we run a hyperparameter sweep over , layer indices 3,4,5, 4,5,6, and 5,6,7, and learning rates . We select runs with the lowest possible WMDP accuracies that don’t completely destroy MMLU accuracy. For Mistral, this is and lr= at layers 5,6,7. We sweep across the same hyperparameters for Llama-3-8B. Llama-3-8B-Instruct uses the best hyperparameters found in the base model sweep. The runs shown in Table 1 have and lr= at layers 4,5,6. All runs had a steering coefficient of 6.5.
RepNoise (Rosati et al., 2024).
Repurposing the authors’ original code, we train RepNoise on Zephyr-7B using the WMDP retain and forget datasets as and respectively. We trained LoRA adapters on top of the original model with rank 64, alpha=16, and dropout=0.05. We first conducted a grid search over the parameters , , and learning rates . As none of the resulting runs significantly decreased accuracy on WMDP MCQ questions without destroying MMLU accuracy, we performed one more grid search over parameters , , and learning rates . The highest-performing run, shown in Table 1, had , , and learning rate . The method was run for one epoch with a batch size of 4.
For Mistral, we run a hyperparameter sweep over , , and learning rates . We selected the run that has the lowest possible WMDP accuracies without destroying MMLU accuracy. This run, shown in Table 1, has the parameters , , lr=.
We run a sweep over the same hyperparameters for Llama-3-8B, and use the best runs from the base model to decide hyperparameters for Llama-3-8B-Instruct. The runs shown in Table 1 had , , lr=.
B.2 Harry Potter Results
RMU (Li et al., 2024).
We train LoRA adapters on top of Llama-2-7B Chat at varying layers, using text from the Harry Potter books (https://huggingface.co/datasets/KaungHtetCho/Harry_Potter_LSTM) as and WikiText as . We sweep across layer indices 3,4,5, 4,5,6, and 5,6,7 with and learning rate . We report numbers for the best run in Table 3, for layers 5,6,7, , learning rate , and batch size 1, trained for one epoch. The Harry Potter dataset used for RMU was not the exact same dataset used for ELM (https://huggingface.co/datasets/mickume/harry_potter_tiny), as performance was much worse for RMU on the latter dataset.
WHP (Eldan & Russinovich, 2023).
We directly download the best Llama-2-7B Chat model from the original authors (https://huggingface.co/microsoft/Llama2-7b-WhoIsHarryPotter).
Appendix C Robustness Evaluation
To evaluate the robustness of ELM against adversarial attacks, we employ the Greedy Coordinate Gradient (GCG) method (Zou et al., 2023), utilizing the standard implementation from GraySwanAI (GraySwanAI, 2024). The GCG attack requires defining an initial prompt, a multi-token target text, and an initialized adversarial suffix. Following the protocol established in Li et al. (2024), we use a 20-token adversarial suffix and derive prompts from the WMDP MCQ datasets. To facilitate open-ended generation, we present only the question component of these prompts, omitting the multiple-choice structure. Our experiments reveal a stark contrast in robustness between ELM models and their base model counterparts. Even after extensive optimization exceeding 5000 iterations, we fail to identify a GCG prompt capable of inducing ELM models to generate content related to erased concepts. This resilience stands in marked contrast to the original models, which succumb to effective attack suffixes within 200 iterations, subsequently producing potentially harmful text.
ELM:
RMU:
RepNoise:
Appendix D Hyperparameter Analysis
To optimize the performance of ELM, we conduct an extensive hyperparameter study, focusing on three key parameters: LoRA rank, erasure strength , and the range of layers to which ELM is applied. Our findings corroborate and extend previous observations in the literature (Meng et al., 2022; Geva et al., 2023). Figure 4a illustrates the impact of layer selection on erasure efficacy.
Consistent with prior work, we observe that applying ELM to earlier layers yields more effective knowledge erasure compared to later layers. Specifically, we identified layers 4-7 of the Zephyr model as the optimal range for achieving a balance between thorough knowledge erasure and preservation of general capabilities.
The interplay between LoRA rank and erasure strength is depicted in Figure 4b. Our analysis reveals that lower values of result in diminished effects on both erasure performance and general benchmark scores. Interestingly, we found no clear trend with respect to LoRA rank, with lower-rank updates performing comparably to higher-rank alternatives. This suggests that ELM can achieve effective erasure with minimal parametric overhead.
Based on these empirical results, we adopted a configuration of rank 4, , and application to layers 4-7 for all subsequent experiments. This configuration strikes a balance between erasure efficacy, computational efficiency, and preservation of general language capabilities.
Appendix E Low-Rank vs Full Finetuning
We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints. In Table 4, we show the performance of ELM on Zephyr-7B when editing with full finetuning and low-rank model editing. Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model. Where as low-rank model editing preserves the specificity while being effective at erasure.
| Method | Innocence () | Specificity () | ||
| Bio | Cyber | MMLU | MT-Bench | |
| Original | 64.4 | 44.3 | 58.5 | 7.3 |
| ELM - Full | 25.4 | 27.1 | 45.2 | 3.4 |
| ELM - LoRA | 29.7 | 27.2 | 56.6 | 7.1 |
Appendix F Conditional Fluency Training
To achieve seamless editing, ELM must generate fluent text even when prompted about erased concepts. The ideal behavior mimics a model that never encountered the concept during pretraining. We implement an additional step to make ELM models acknowledge the concept while suggesting a topic change, although this behavior remains configurable through prompt engineering.
Our training procedure extends the erasure objective from Equation 10. For each prompt from the harmful dataset, we generate new tokens using the erasure objective. Importantly, we do not consider these newly generated tokens as harmful context for subsequent generations, but rather use them for positive conditioning. This approach allows the model to continue generating fluently while reducing the likelihood of discussing the erased concept. Through this process, the model learns to maintain fluency while decreasing the probability of elaborating on the queried concept. Inspired by Qi et al. (2024), we incorporate an additional consistency mechanism. We append a standard response to the initial prompt, such as a paraphrased version of: “This is a harmful concept. Let’s change the topic to something more fun and interesting:” We then initiate the generation process from this augmented prompt. This technique ensures consistent model behavior when encountering erased concepts. The final training step involves generating the complete response, including the initial prompt, consistency prompt, and letting the model generate new tokens. We then pass this entire sequence through the ELM model. Crucially, we fine-tune only the parameters responsible for generating the new tokens. This targeted approach ensures that we preserve the model’s general knowledge while specifically adapting its behavior for erased concepts.
Appendix G Progression of ELM Training
We evaluate the ELM intermediate checkpoints to observe the training dynamics of the method in Figure 5. We find that ELM suddenly drops the knowledge of the erased concept, halfway down the training and continues to slowly erase the rest of the traces. Bio-threat knowledge takes more time to be erased from the model - which could be directly proportional to the initial amount of prior knowledge.
Appendix H Adversarial Attacks
We also attack ELM with BEAST (Sadasivan et al., 2024), a fast adversarial prompt based attack on LLMs. BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response. We find that BEAST is unable to extract erased information from ELM:
Additionally - finetuning attack where we train ELM model autoregressively on the original forget dataset. We find that the resulting attacked model brings back the knowledge slightly (Bio: 29.7% to 42.2%; Cyber: 27.2% to 29.4%) but not to the original level of 64.4% Bio and 44.3% Cyber. ELM models can be retrained to bring back erased knowledge, but it is harder.